OpenStudy (anonymous):

Find the complete factored form of -25a^6 + 10a^5 + 45a^4

OpenStudy (whpalmer4):

Are there any common factors to the coefficients? If so, factor them out by using the distributive property in reverse: ab + ac + ad = a(b+c+d) Are there any common factors to the rest of the terms? Again, if so, factor them out. It may be helpful as a learning exercise to write it out as its factors. For example, \[3x^2 + 6x = 3*x*x + 2*3*x = 3*(x*x + 2*x) = 3*x*(x+2) = 3x(x+2)\]

OpenStudy (whpalmer4):

Last line got cut off, it should end with \[= 3x(x+2)\]

OpenStudy (anonymous):

I think I got it, thanks!

OpenStudy (whpalmer4):

What's your answer?

OpenStudy (anonymous):

5a^4 (-5a^2 + 2a + 9)

OpenStudy (anonymous):

I don't think that's complete though...

OpenStudy (whpalmer4):

Okay, that's correct so far, but your instinct that we need to factor the trinomial \(-5a^2 + 2a + 9\) is a good one. Do you know to do that?

OpenStudy (whpalmer4):

As it turns out, while the instinct is good, in this case the trinomial can't be further factored into a product of terms such as \[(-5a + 7)(a + 1)\] (which would give us \(-5a^2-5a+7a +7 = -5a^2+2a+7\) if expanded—close, but no cigar!)